OpenStudy (anonymous):

Solve (1/8)r = 1/4 MEDAL WILL BE GIVEN!(;

OpenStudy (anonymous):

What do you get when you multiply both sides of your equation by 8? :-)

OpenStudy (anonymous):

I cant think properly right now lol thats why Im asking. @ChristopherToni

OpenStudy (anonymous):

You can take the reciprocal of 1/8 and multiply it to both sides to isolate your r variable.

OpenStudy (anonymous):

Multiplying both sides by 8, we see that \(8 \cdot \dfrac{1}{8}r = 8\cdot \dfrac{1}{4} \implies \dfrac{8}{8}r = \dfrac{8}{4}\). How does this simplify?

OpenStudy (zzr0ck3r):

you shuold do homework when you are thinking properly. Or ask some questions. You are basically asking the same question over and over. If you learn how to do it once, you will no longer need to ask these quesetions.

OpenStudy (zzr0ck3r):

questions*

OpenStudy (anonymous):

I cant wait till I am thinking properly @zzr0ck3r I have to have this done as soon as possible.

OpenStudy (anonymous):

anyways thanks for your helps guys